find the differential equation:

2x + 6 y y' =(x^2+3y^2/y)y'


1
Expert's answer
2021-05-24T16:24:56-0400

Solution.


"2x+6yy'=(x^2+3y^2)\/y\u2022y'"

Let be "y=x\u2022v(x)," then

"y'=xv'(x)+v."

We will have


"2x+6x(xv'+v)•v=\frac{(x^2+3x^2v^2)(xv'+v)}{xv}"

"2x+6x(xv'+v)•v=\frac{x(1+3v^2)(xv'+v)}{v}"

"2xv+6x^2v^2v'+6xv^3=3x^2v^2v'+x^2v'+3xv^3+xv"

"xv+3x^2v^2v'+3xv^3=x^2v'"

"x(3v^2-1)v'=-3v^3-v"

"\frac{3v^2-1}{-3v^3-v}v'=\frac{1}{x}"


"\int\frac{3v^2-1}{-3v^3-v}dv=\int\frac{1}{x}dx"

"-\int(\frac{6v}{3v^2+1}-\frac{1}{v})dv=\int\frac{1}{x}dx"

"-\ln (3v^2+1)+\ln v=\ln x+C,"

where C is some constant.


"v=x(3v^2+1)e^C"

"y=(3y^2+x^2)e^C"

"-3e^Cy^2+y-x^2e^C=0"

"y=\frac{1}{6}(e^C\pm\sqrt{e^{2C}-12x^2})."